CEDA News

Viewing posts from October, 2007

BADC Server Issue - Number of datasets affected

Posted by: CEDA_site Admin in Other 9 years, 5 months ago
Following a technical issue with one of our servers the following datasets are not accessible at present: caesar covex dabex fluxex iasi icepic land-emiss micromix nu-wave raptor rico t-rex visurb Meris l1b data We hope to have the Meris data back shortly and the other datasets back in a few days. Further news items will be posted as progress is made. We apologise for any inconvenience caused.

BADC Helpdesk 25th October 2007

read more / 0 comments

HadCM3-control data re-instated into archive.

Posted by: CEDA_site Admin in Met Office data 9 years, 5 months ago
HadCM3-control data have been successfully restored from backup and is fully accessible once more from the BADC. Further news items regarding the server issues will be posted as progress is made. Once again we apologise for any inconvenience caused.

BADC Helpdesk 25th October 2007

read more / 0 comments

BADC Server Issue - Hadcm3-control and UM data affected

Posted by: CEDA_site Admin in Other 9 years, 5 months ago
Following a technical issue with one of our servers the entire HADCM3-control data and approximately 1/6th of the UM data are currently unavailable. We hope to have the HadCM3-control data back early next week and will endeavour to have the UM data re-instated into the archive as soon as possible. Further news items will be posted as progress is made. We apologise for any inconvenience caused.

BADC Helpdesk 17th October 2007

read more / 0 comments

BADC FTP server Re-configuration - COMPLETED

Posted by: CEDA_site Admin in Other 9 years, 5 months ago
Re-configuration of the FTP server has completed successfully. We hope that the short outages did not cause any inconvenience. Please let us know if you encounter any problems.

BADC Helpdesk 4th October 2007

read more / 0 comments

BADC FTP server Re-configuration

Posted by: CEDA_site Admin in Other 9 years, 5 months ago
Tomorrow (3rd October 2007) the BADC ftp server is being re-configured. The changes should not take long but there could be some interruption to the service throughout the day. A further news item will be issued once the work is complete.

Apologies for any inconvenience this essential work may cause.

read more / 0 comments

CEDA News

Recent Posts

Archived Posts

[ months ] [ categories ] [ tags ] [ authors ] [ feeds ]

Months

2017

2016

2015

2014

2013

2012

2011

2010

2009

2008

2007

2006

2005

2004

2003

2002

2001

2000

1996

Categories

Tags

Authors

Feeds

This website and others run by CEDA uses cookies. By continuing to use this website you are agreeing to our use of cookies.

OK | Find out more